# A contramodule generalization of Neeman’s flat and projective module theorem

### v3

@@ Introduction
We also consider the exact category of cotorsion $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$. According to [36, Theorem 7.17 and Corollary 7.21], for any associative ring $R$, the inclusion of exact/abelian categories $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ (where $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}}$ is the full subcategory of cotorsion left $R$-modules) induced equivalences of their derived and contraderived categories $\mathsf{D}(R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}})\simeq\mathsf{D}(R{\operatorname{\mathsf{--Mod}}})$ and $\mathsf{D}^{\mathsf{bctr}}(R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}})\simeq\mathsf{D}^{\mathsf{bctr}}(R{\operatorname{\mathsf{--Mod}}})$. For a complete, separated topological ring $\mathfrak{R}$ with a countable base of neighborhoods of zero consisting of two-sided ideals, it was shown in [34, Corollary 12.8] that the inclusion of exact/abelian categories $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ induces an equivalence of the derived categories $\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})\simeq\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$. In this paper we show that the same inclusion induces an equivalence of the contraderived categories $\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$.
@@ 2. Contraderived Categories
In the stated generality, a proof can be found in [38, Lemma 4.5]; see also [40, Proposition 1.3]. ∎
@@
These observations seem to go back to [28, Section 3.5]. Part (b) is obvious. In part (a), the point is that, for every complex of projective objects $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}\in\mathsf{B}$, the short seqeuence of complexes of abelian groups $0\longrightarrow\operatorname{Hom}_{\mathsf{B}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(P^{\text{\smaller\smaller$\scriptstyle\bullet$}},K^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}_{\mathsf{B}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(P^{\text{\smaller\smaller$\scriptstyle\bullet$}},L^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}_{\mathsf{B}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(P^{\text{\smaller\smaller$\scriptstyle\bullet$}},M^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow 0$ is exact. The total complex of any short exact sequence of complexes of abelian groups is exact. For a generalization, see Lemma 7.5 below. ∎
@@ 3. Topological Rings
Now let $\mathfrak{R}$ be a complete, separated topological ring with a base of neighborhoods of zero formed by open right ideals. Then the functor $\mathfrak{R}[[{-}]]\colon\mathsf{Sets}\longrightarrow\mathsf{Sets}$ has a natural structure of a monad on the category of sets (in the sense of [25, Chapter VI]). For any set $X$, the monad unit $\epsilon_{\mathfrak{R},X}\colon X\longrightarrow\mathfrak{R}[[X]]$ is the map taking every element $x\in X$ to the formal linear combination $\sum_{y\in X}r_{y}y$ with $r_{x}=1$ and $r_{y}=0$ for $y\neq x$. The monad multiplication $\phi_{\mathfrak{R},X}\colon\mathfrak{R}[[\mathfrak{R}[[X]]]]\longrightarrow\mathfrak{R}[[X]]$ is the “opening of parentheses” map assigning a formal linear combination to a formal linear combination of formal linear combinations. In order to open the parentheses, one needs to compute pairwise products of elements of the ring $\mathfrak{R}$ and certain infinite direct sums in $\mathfrak{R}$, which are understood as the limits of finite partial sums in the topology of $\mathfrak{R}$. The conditions imposed on $\mathfrak{R}$ (including, in particular, the requirement of a topology base of open right ideals) guarantee the convergence.
In the general category-theoretic terminology, one usually speaks about “algebras over monads”, but in our context we prefer to call them modules over monads. A left $\mathfrak{R}$-contramodule is defined as a module over the monad $X\longmapsto\mathfrak{R}[[X]]$ on the category of sets. In other words, a left $\mathfrak{R}$-contramodule $\mathfrak{P}$ is a set endowed with a left contraaction map $\pi_{\mathfrak{P}}\colon\mathfrak{R}[[\mathfrak{P}]]\longrightarrow\mathfrak{P}$ satisfying the following contraassociativity and contraunitality axioms. Firstly, the map $\mathfrak{R}[[\pi_{\mathfrak{P}}]]\colon\mathfrak{R}[[\mathfrak{R}[[\mathfrak{P}]]]]\longrightarrow\mathfrak{R}[[\mathfrak{P}]]$ induced by the map $\pi_{\mathfrak{P}}$ and the “opening of parentheses” map $\phi_{\mathfrak{R},\mathfrak{P}}\colon\mathfrak{R}[[\mathfrak{R}[[\mathfrak{P}]]]]\longrightarrow\mathfrak{R}[[\mathfrak{P}]]$ must have equal compositions with the contraaction map $\pi_{\mathfrak{P}}$,
@@
A right $\mathfrak{R}$-module $\mathcal{N}$ is said to be discrete if the action map $\mathcal{N}\times\mathfrak{R}\longrightarrow\mathcal{N}$ is continuous in the given topology of $\mathfrak{R}$ and the discrete topology of $\mathcal{N}$. The full subscategory of discrete right $\mathfrak{R}$-modules ${\operatorname{\mathsf{Discr--}}}\mathfrak{R}\subset{\operatorname{\mathsf{Mod--}}}\mathfrak{R}$ is closed under subobjects, quotients, and infinite direct sums in the category of right $\mathfrak{R}$-modules ${\operatorname{\mathsf{Mod--}}}\mathfrak{R}$. The category ${\operatorname{\mathsf{Discr--}}}\mathfrak{R}$ is a Grothendieck abelian category.
@@ 4. Countably Presentable Contramodules
Let $\mathfrak{R}$ be a complete, separated topological ring, and let $\mathfrak{R}\supset\mathfrak{I}_{1}\supset\mathfrak{I}_{2}\supset\mathfrak{I}_{3}\supset\dotsb$ be a countable base of neighborhoods of zero in $\mathfrak{R}$ consisting of open two-sided ideals $\mathfrak{I}_{n}$,  $n\geq 1$. Denote by $R_{n}=\mathfrak{R}/\mathfrak{I}_{n}$ the corresponding discrete quotient rings. Then the rule assigning to a flat left $\mathfrak{R}$-contramodule $\mathfrak{F}$ the collection of $R_{n}$-modules $F_{n}=\mathfrak{F}/(\mathfrak{I}_{n}\rightthreetimes\mathfrak{F})$ is an equivalence between the category of flat left $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ and the category of following sets of data:
@@ 5. Acyclic Complexes of Flat Contramodules are Contraacyclic
It remains to recall that the homological dimension of the exact category of countably presentable flat $\mathfrak{R}$-contramodules does not exceed $1$ by Proposition 4.7. By Lemma 2.7, it follows that the complex $\mathfrak{H}$ is absolutely acyclic in the exact category of countably presentable flat $\mathfrak{R}$-contramodules, and consequently also in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ and in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. By Lemma 2.5(a), any morphism of complexes $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero. ∎

### v4

@@ Introduction
We also consider the exact category of cotorsion $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}$. According to [36, Theorem 7.17 and Corollary 7.21], for any associative ring $R$, the inclusion of exact/abelian categories $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ (where $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}}$ is the full subcategory of cotorsion left $R$-modules) induces equivalences of their derived and contraderived categories $\mathsf{D}(R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}})\simeq\mathsf{D}(R{\operatorname{\mathsf{--Mod}}})$ and $\mathsf{D}^{\mathsf{bctr}}(R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}})\simeq\mathsf{D}^{\mathsf{bctr}}(R{\operatorname{\mathsf{--Mod}}})$. For a complete, separated topological ring $\mathfrak{R}$ with a countable base of neighborhoods of zero consisting of two-sided ideals, it was shown in [34, Corollary 12.8] that the inclusion of exact/abelian categories $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}\longrightarrow\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ induces an equivalence of the derived categories $\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})\simeq\mathsf{D}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$. In this paper we show that the same inclusion induces an equivalence of the contraderived categories $\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}})\simeq\mathsf{D}^{\mathsf{bctr}}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$.
@@ 2. Contraderived Categories
In the stated generality, a proof can be found in [38, Lemma 4.5]; see also [40, Proposition 1.3] and [42, Lemma 7.5]. ∎
@@
These observations seem to go back to [28, Section 3.5] and [22, Proposition 4.3]. Part (b) is obvious. In part (a), the point is that, for every complex of projective objects $P^{\text{\smaller\smaller$\scriptstyle\bullet$}}\in\mathsf{B}$, the short sequence of complexes of abelian groups $0\longrightarrow\operatorname{Hom}_{\mathsf{B}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(P^{\text{\smaller\smaller$\scriptstyle\bullet$}},K^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}_{\mathsf{B}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(P^{\text{\smaller\smaller$\scriptstyle\bullet$}},L^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}_{\mathsf{B}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(P^{\text{\smaller\smaller$\scriptstyle\bullet$}},M^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow 0$ is exact. The total complex of any short exact sequence of complexes of abelian groups is exact. For a generalization, see Lemma 7.5 below. ∎
@@ 3. Topological Rings
Now let $\mathfrak{R}$ be a complete, separated topological ring with a base of neighborhoods of zero formed by open right ideals. Then the functor $\mathfrak{R}[[{-}]]\colon\mathsf{Sets}\longrightarrow\mathsf{Sets}$ has a natural structure of a monad on the category of sets (in the sense of [25, Chapter VI]). For any set $X$, the monad unit $\epsilon_{\mathfrak{R},X}\colon X\longrightarrow\mathfrak{R}[[X]]$ is the map taking every element $x\in X$ to the formal linear combination $\sum_{y\in X}r_{y}y$ with $r_{x}=1$ and $r_{y}=0$ for $y\neq x$. The monad multiplication $\phi_{\mathfrak{R},X}\colon\mathfrak{R}[[\mathfrak{R}[[X]]]]\longrightarrow\mathfrak{R}[[X]]$ is the “opening of parentheses” map assigning a formal linear combination to a formal linear combination of formal linear combinations. In order to open the parentheses, one needs to compute pairwise products of elements of the ring $\mathfrak{R}$ and certain infinite sums of elements of $\mathfrak{R}$, which are understood as the limits of finite partial sums in the topology of $\mathfrak{R}$. The conditions imposed on $\mathfrak{R}$ (including, in particular, the requirement of a topology base of open right ideals) guarantee the convergence.
In the general category-theoretic terminology, one usually speaks about “algebras over monads”, but in our context we prefer to call them modules over monads. A left $\mathfrak{R}$-contramodule is defined as a module over the monad $X\longmapsto\mathfrak{R}[[X]]$ on the category of sets. In other words, a left $\mathfrak{R}$-contramodule $\mathfrak{P}$ is a set endowed with a left contraaction map $\pi_{\mathfrak{P}}\colon\mathfrak{R}[[\mathfrak{P}]]\longrightarrow\mathfrak{P}$ satisfying the following contraassociativity and contraunitality axioms. Firstly, the “opening of parentheses” map $\phi_{\mathfrak{R},\mathfrak{P}}\colon\mathfrak{R}[[\mathfrak{R}[[\mathfrak{P}]]]]\longrightarrow\mathfrak{R}[[\mathfrak{P}]]$ and the map $\mathfrak{R}[[\pi_{\mathfrak{P}}]]\colon\mathfrak{R}[[\mathfrak{R}[[\mathfrak{P}]]]]\longrightarrow\mathfrak{R}[[\mathfrak{P}]]$ induced by the map $\pi_{\mathfrak{P}}$ must have equal compositions with the contraaction map $\pi_{\mathfrak{P}}$,
@@
A right $\mathfrak{R}$-module $\mathcal{N}$ is said to be discrete if the action map $\mathcal{N}\times\mathfrak{R}\longrightarrow\mathcal{N}$ is continuous in the given topology of $\mathfrak{R}$ and the discrete topology of $\mathcal{N}$. The full subcategory of discrete right $\mathfrak{R}$-modules ${\operatorname{\mathsf{Discr--}}}\mathfrak{R}\subset{\operatorname{\mathsf{Mod--}}}\mathfrak{R}$ is closed under subobjects, quotients, and infinite direct sums in the category of right $\mathfrak{R}$-modules ${\operatorname{\mathsf{Mod--}}}\mathfrak{R}$. The category ${\operatorname{\mathsf{Discr--}}}\mathfrak{R}$ is a Grothendieck abelian category.
@@ 4. Countably Presentable Contramodules
Let $\mathfrak{R}$ be a complete, separated topological ring, and let $\mathfrak{R}\supset\mathfrak{I}_{1}\supset\mathfrak{I}_{2}\supset\mathfrak{I}_{3}\supset\dotsb$ be a countable base of neighborhoods of zero in $\mathfrak{R}$ consisting of open two-sided ideals $\mathfrak{I}_{n}$,  $n\geq 1$. Denote by $R_{n}=\mathfrak{R}/\mathfrak{I}_{n}$ the corresponding discrete quotient rings. Then the rule assigning to a flat left $\mathfrak{R}$-contramodule $\mathfrak{F}$ the collection of $R_{n}$-modules $F_{n}=\mathfrak{F}/(\mathfrak{I}_{n}\rightthreetimes\mathfrak{F})$ is an equivalence between the category of flat left $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ and the category of the following sets of data:
@@ 5. Acyclic Complexes of Flat Contramodules are Contraacyclic
It remains to recall that the homological dimension of the exact category of countably presentable flat $\mathfrak{R}$-contramodules does not exceed $1$ by Proposition 4.7. By Lemma 2.7, it follows that the complex $\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is absolutely acyclic in the exact category of countably presentable flat $\mathfrak{R}$-contramodules, and consequently also in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ and in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. By Lemma 2.5(a), any morphism of complexes $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero. ∎